OpenStudy (anonymous):

Part 1 − Find the vertex, axis of symmetry, domain, and range of the graph of y = −3x2 − 3x + 4. Show all work for full credit. Part 2 − Using complete sentences, explain how you can determine the axis of symmetry, the domain, and range without graphing y = −3x2 − 3x + 4.

OpenStudy (anonymous):

Do you know how to find the vertex?

OpenStudy (anonymous):

First of all the coefficient of \[x ^{2}\] is negatif (i.e -3)

OpenStudy (anonymous):

so the parabola has an maximum

OpenStudy (anonymous):

you have to find the roots using the quadratic formula

OpenStudy (anonymous):

\[x _{1,2}=\frac{ -b \pm \sqrt{b ^{2}-4ac} }{ 2a}\]

OpenStudy (anonymous):

x1=\[\frac{ 3+\sqrt{9-4(-3)4} }{ 2(-3) }\]

OpenStudy (anonymous):

x1=-1.75

OpenStudy (anonymous):

x2=0.75

OpenStudy (anonymous):